b) Uwezo Ltd intends to issue additional funds as follows: – 100,000 ordinary shares of shs 20 each

50,000 10% preference shares of shs 10 each

40,000 10% Debentures; @ shs 100 The market values are;

Ordinary shares shs 30, floatation fees of shs 2 per share Preference shares sh. 12, floatation fees of sh. 1 per share Debenture shs 90, and floatation fee shs 5 per debenture

 


Required:

Calculate weighted average cost of capital of Uwezo Ltd.

1
Expert's answer
2021-07-07T13:48:35-0400

"WACC=(E\/V)*Re+(D\/V)*Rd(1-Tc)+(P\/V)*Rp"

Ordinary share capital= "(30-2)*100000=2,800,000"

Preference share capital = "(12-1)*50000=550,000"

Value of Debenture ="(90-5)*40000 =3,400,000"

Total value of debt and equity= "2,800,000+550,000+3,400,000 = 6,750,000"

"WACC=2,800,000\/6,750,000*10%+550,000\/6,750,000*10%+3,400,000\/6,750,000*10%""+550,000\/6,750,000*10%""+3,400,000\/6,750,000*10%"

"WACC=4.15+0.82+5.04"

"WACC=10.01%"%